OpenStudy (anonymous):

calculated the derivative with respect to x: 4x^4 + 3y^5 =7x + 9y ?

OpenStudy (hoblos):

16x^3 + 15y^4y' = 7 +9y'

OpenStudy (anonymous):

then 16x^3-7 = 9y'-15y^4y' ?

OpenStudy (hoblos):

yup

OpenStudy (anonymous):

so i got y' = (16x^3-7) all over 9-15y^4 ?

OpenStudy (hoblos):

yup

OpenStudy (anonymous):

thanks!

OpenStudy (hoblos):

welcome :D